) In a study of TV viewing habits, it is desired to estimate the average number of hours that learners spend




watching per week. If it is reasonable to assume that




  3.2




hours, how large a sample is needed so that it




will be possible to assert with 95% confidence that the sample mean is off by less than 20 minutes? �

1
Expert's answer
2022-01-25T11:15:16-0500

"Given~\sigma=4.8hours\\
|X-\mu|<1.96\times\frac{4.8}{\sqrt{n}}\\
1.96\times\frac{4.8}{\sqrt{n}}>\frac{1}{3}\\
\frac{9.408}{\sqrt{n}}>\frac{1}{3}\\
\frac{1}{\sqrt{n}}>\frac{1}{28.224}\\
n=28.224^2\\
n=796.59"